OpenStudy (vianne):

The simple interest on an investment of $4000 for one year is $140. What is the interest rate?

OpenStudy (texaschic101):

I = PRT....however, since we are looking for R, we can change the equation to this : I/PT = R I = 140 P = 4000 T = 1 now we sub I/PT = R 140/(4000)(1) = R 140/4000 = R 0.035 = R (change to a percent) 3.5% = R check... I = PRT 140 = (4000)(.035)(1) 140 = 140 (correct) Interest rate is 3.5%